OpenStudy (beth12345):

how to find the value of 'X' when x^2-8x=8?

OpenStudy (mimi_x3):

\[x^2-8x-8 = 0 \] Use the quadratic formula: \[\frac{-(b)\pm\sqrt{(b)^{2}-4(a)(c)}}{2(a)} \] \[a = 1 , b=-8 , c = -8\] And sub it in.

OpenStudy (beth12345):

ok thanx

OpenStudy (anonymous):

use the quadratic formula. \[(-b \pm \sqrt{b ^{2} -4ac}) /2a \]

OpenStudy (anonymous):

x^2-8x=8 x(x-8)=8 or x=2 and x-8=4 x=2, x=12

OpenStudy (beth12345):

thank you, i never thouht of it that way but i can see how it works :)

OpenStudy (anonymous):

but this solution is wrong

OpenStudy (beth12345):

?

OpenStudy (anonymous):

x=2, x=12 is a wrong solution

OpenStudy (mimi_x3):

Completing the square would work as well. But using the quadratic formula is easier.

OpenStudy (beth12345):

i thought it would be like (2)[(12)-8]=8?

OpenStudy (beth12345):

oh ok, mimi, i did the formula and got 4+or-2root6, is that right?

OpenStudy (anonymous):

Ravi, in case you were wondering why it's wrong it's because you must factorise the equation equal to 0. 8 has an infinite number of non-integer factors, so you can't solve. 0, on the other hand, must have the value of a factor as 0.

OpenStudy (beth12345):

thanks samarilli, i see now why it's not the same answer as when i use the formula

OpenStudy (mimi_x3):

Well, i got \[\frac{-8\pm\sqrt{96}}{2} \]

OpenStudy (anonymous):

thanks Simarilli